# Giovanni Compiani Ilya Morozov Stephan Seiler

### v2

@@
Demand Estimation with Text and Image Datathanks: We thank Steve Berry, Alex Burnap, Matt Gentzkow, Joonhwi Joo, Philip Haile, Yufeng Huang, Ayush Kanodia, Carl Mela, Sanjog Misra, Olivia Natan, Matt Osborne, Katja Seim, Adam Smith, Yuyan Wang, Kevin Williams and seminar participants at Bocconi University, Chicago Booth, London Business School, Stockholm School of Economics, University of Cologne, Nova SBE, Marketing Science Conference 2022, SICS Conference 2024, AI/ML Yale Conference, and SED Conference for helpful comments. We also thank Celina Park, Tanner Parsons, James Ryan, Janani Sekar, Andrew Sharng, Franklin She, Adam Sy, and Vitalii Tubdenov for outstanding research assistance. Contact information: Giovanni Compiani (gio1compiani@gmail.com), Ilya Morozov (ilya.morozov@kellogg.northwestern.edu), Stephan Seiler (stephan.a.seiler@gmail.com). Giovanni Compiani gratefully acknowledges support from the Center for Applied AI at Booth.
@@
We propose a demand estimation method that leverages unstructured text and image data to infer substitution patterns. Using pre-trained deep learning models, we extract embeddings from product images and textual descriptions and incorporate them into a random coefficients logit model. This approach enables researchers to estimate demand even when they lack data on product attributes or when consumers value hard-to-quantify attributes, such as visual design or functional benefits. Using data from a choice experiment, we show that our approach outperforms standard attribute-based models in counterfactual predictions of consumers’ second choices. We also apply it across 40 product categories on Amazon.com and consistently find that text and image data help identify close substitutes within each category.
@@ 1 Introduction
In this paper, we show how researchers can incorporate text and image data in demand estimation to recover substitution patterns. Our approach uses product images, titles, descriptions, and customer reviews—unstructured data that are widely available even in markets where collecting product attributes is challenging. Using pre-trained deep learning models, we extract low-dimensional features from images and texts and include them in a standard logit demand model, interacting them with random coefficients to capture substitution patterns. This approach enables researchers to incorporate hard-to-quantify product attributes, such as visual design from images and functional benefits from text, while also circumventing the need to subjectively choose which observed attributes to collect.
To validate our approach, we run an experiment that elicits consumers’ first and second choices. We estimate our model on first choices and show that it outperforms standard attribute-based demand models at counterfactually predicting second choices—a key empirical measure of substitution. Our findings suggest that researchers can use text and images instead of traditional attribute data and that these unstructured data may even capture substitution beyond what is reflected in observed product attributes. Lastly, we apply our approach to e-commerce data and consistently find that text and images effectively identify close substitutes across many product categories.
Proposed Approach and Validation
@@
To extract information from unstructured data, we use pre-trained deep learning models, transforming images and texts into vector representations—embeddings. We then apply principal component analysis (PCA) to these embeddings to further reduce dimensionality and capture the main dimensions of product differentiation. This process generates separate embeddings and principal components for each data source: product images, titles, descriptions, and customer reviews. We incorporate these principal components into a mixed logit demand model (Berry et al., 1995; Nevo, 2001), interacting them with random coefficients similar to how researchers usually treat observed product attributes in demand estimation. To identify the principal components that best capture substitution in the data, we select the specification that minimizes the Akaike Information Criterion (AIC).
Just as standard demand models capture substitution through the similarity of observed attributes, our approach captures it by looking at how close products are in the space of texts and images. This proximity will reflect substitution if consumers consider how a product looks and how it is described by sellers and other consumers. Alternatively, visual and textual descriptions may serve as proxies for product features and functional benefits. Take tablets, for example: product titles may reveal brand and screen size, descriptions may highlight whether the tablet is suitable for drawing or gaming, reviews may mention that a tablet is durable and child-friendly, and photos may showcase design features like color and casing style. By extracting embeddings from texts and images, we identify these dimensions of differentiation.
To evaluate our approach, we run an experiment where participants choose a book from a list of options. We randomize both the prices and rankings of books to generate the variation needed to identify substitution patterns (Berry and Haile, 2014). We elicit participants’ first and second choices, estimate each demand model using only first choices, and then counterfactually predict second choices. Because second choices reveal which books consumers substitute to when their first choice is unavailable, predicting them well requires a demand model that accurately captures substitution patterns.
We find that incorporating either text or images in demand estimation improves both in-sample fit and counterfactual predictions relative to the plain logit model without random coefficients. Images capture substitution because book covers visually convey genre and distinguish between fiction and non-fiction. Text performs even better, especially when we use more advanced models and richer text data. This is likely because descriptions and reviews capture nuanced details about book plots, helping identify similar books even within the same genre. Crucially, our best-fitting specification—based on customer reviews—outperforms standard attribute-based demand models in counterfactual predictions. These findings suggest that text and images can replace standard attribute data and may even capture substitution patterns beyond what is reflected in observed attributes.
Beyond improving fit and counterfactual second-choice predictions, our approach accurately identifies the closest substitutes among products. We show that for some books our method correctly recovers the five closest substitutes, whereas alternative models misidentify them. Additionally, we simulate hypothetical mergers of book publishers and find that demand models diverge in their predictions of expected price increases. This divergence has significant implications for downstream decisions. For instance, if our model captures substitution accurately, an antitrust agency relying on a plain logit model or an attribute-based mixed logit model will incorrectly approve mergers that should be challenged and block mergers that should be approved.
Lastly, we apply our approach to 40 product categories on Amazon.com, including groceries, pet food, office supplies, beauty products, electronics, video games, and clothing. We combine data on online purchases from the Comscore Web Behavior panel with product images, titles, descriptions, and reviews collected from Amazon’s product detail pages. Across all categories, we consistently find that text and images contain useful information about substitution patterns. This additional information helps identify close substitutes within each category, leading to substantial deviations from restrictive substitution patterns in the plain logit model. These results confirm that our approach performs consistently well across a wide range of products, even in observational data that has less price variation than our experiment.
Contributions
Our main innovation is to show how researchers can incorporate unstructured text and image data in demand estimation to better recover substitution patterns. Standard demand estimation methods assume a pre-determined nesting structure (McFadden, 1974a), account for preference heterogeneity over observed product attributes (Berry et al., 1995; McFadden and Train, 2000; Berry et al., 2004), or parameterize utility covariances based on distances between products in attribute space (Bresnahan et al., 1997; Pinkse and Slade, 2004; Dotson et al., 2018; Chintagunta and Li, 2024). [^2] The idea of treating products as bundles of characteristics dates back to Lancaster (1966) and McFadden (1974b). Building on this body of work, we show how researchers can incorporate product images and textual product descriptions into standard demand models. [^3] In concurrent work, Sisodia et al. (2024) extract interpretable product attributes from images and use them as inputs to a conjoint analysis.
@@
Given the prevalence of unstructured data, we see our approach as a valuable addition to the toolbox of empirical researchers. To make it easier for others to apply our approach, we provide a publicly available Python package, DeepLogit, alongside this paper. [^4] The package is available on PyPI and in the public GitHub repository: github.com/deep-logit-demand/deeplogit. Our package heavily borrows from the existing package Xlogit for GPU-accelerated estimation of mixed logit models, developed by Arteaga et al. (2022).
Extracting information from images can be especially valuable in markets where visual design and style matter, such as shoes, clothing, luxury watches, handbags, and art. Similarly, textual descriptions and reviews can capture hard-to-quantify attributes, such as network reliability and customer service for telecom services or usability, customizability, and developer support for software products. Even when product attributes are available, including information from texts and images may still improve demand estimates by capturing additional dimensions of product differentiation.
A natural consequence of using text and images is that our approach scales well across product categories. In contrast, capturing substitution through product attributes across many categories can be challenging. For example, Döpper et al. (2024) remark that doing so “would be difficult to implement at scale because it would require category by-category assessments about which characteristics are appropriate to include and whether or not relevant data are available.” Our approach circumvents the need to collect distinct attributes for each category, making it easier to estimate demand across many categories. This scalability may also be valuable for retailers and platforms seeking to estimate demand across diverse categories and optimize prices at scale.
We validate our approach using both experimental and observational data. Our experiment design lets us directly measure second-choice diversion ratios, allowing us to assess how well an estimated demand model predicts them counterfactually. Diversion ratios reflect substitutability and thus the intensity of price competition between two products. They are crucial for antitrust authorities trying to predict price increases from horizontal mergers (Shapiro, 1995; Conlon and Mortimer, 2021; Conlon et al., 2022), and for managers optimizing assortments and prices. We show that incorporating texts and images into demand models improves diversion ratio estimates, highlighting the value of our approach. Additionally, we use observational data to show that our method performs well across various product categories. We consider this extensive validation with both experimental and observational data an important contribution of the paper.
@@
Several other papers have proposed leveraging text and image data in demand estimation. [^5] Netzer et al. (2012) use data on the co-occurrence of products mentioned in online discussion forums to generate a visual representation of competing products, but they do not incorporate these measures into demand estimation. Quan and Williams (2019) incorporate product image embeddings as shifters of mean utilities in a demand model, showing their approach produces reasonable price elasticity estimates. Similarly, Lee (2024) employs large-language models to predict utility intercepts of new products based on their textual descriptions. These methods use texts or images to predict mean utilities but are not designed to estimate utility covariances, which play a crucial role in shaping substitution. [^6] For example, Lee (2024) inverts a plain logit model without random coefficients to extract product fixed effects for the prediction part of his analysis. Thus, his model restricts substitution patterns due to the IIA property. In contrast, our approach uses the same types of unstructured data to model utility covariances. This enables us to capture flexible substitution patterns—a key input for many empirical applications of interest.
@@
A few papers propose estimating substitution patterns using survey data, a possible alternative to our use of unstructured data. Dotson et al. (2019) have survey participants rate each product image and incorporate rating correlations as utility correlation shifters in their demand model. Similarly, Magnolfi et al. (2022) ask survey participants to rank products relative to each other (e.g., “Product A is closer to B than to C”). They generate low-dimensional product embeddings from these rankings and use them as attributes in a random coefficients logit model. Like our paper, Magnolfi et al.’s approach provides a way to crowdsource product attributes when collecting such data is challenging. While our approach can complement these survey-based methods, it has the advantage of using only widely available text and image data, eliminating the need to conduct costly category-specific surveys. [^7] Dew (2024) also uses survey data to estimate consumer preferences for text and image embeddings. However, his focus is on predicting individual choices rather than estimating demand and modeling substitution.
@@
Our approach is well-suited for several empirical applications. Researchers can use it to address a wide range of economics and marketing questions that require accurate and flexible estimates of product substitution. This includes analyzing how horizontal mergers (Nevo, 2000; Federal Trade Commission, 2022), new product launches (Hausman, 1994; Petrin, 2002), corrective taxes (Allcott et al., 2019; Seiler et al., 2021), and trade restrictions (Goldberg, 1995; Berry et al., 1999) influence consumers’ choices and welfare through changing assortments and prices. Additionally, marketers can apply this approach to help multi-product firms optimize prices and promotions (Hoch et al., 1995; Hitsch et al., 2021).
@@ 2 Proposed Approach
Our approach involves three steps: (1) extracting embeddings from images and texts, (2) reducing the dimensionality of these embeddings using PCA, and (3) including the resulting principal components in a random coefficients logit model. In this section, we focus on the first two steps, as the mixed logit model is standard and does not require a detailed explanation. Section 3.2 describes the exact model specification used in our empirical applications.
2.1 Extracting Embeddings from Texts and Images
We extract embeddings from texts and images using pre-trained deep-learning models. This enables us to leverage models trained on large, general-purpose datasets, which can efficiently capture differences in product images and descriptions. Further, using pre-trained models reduces the computational burden of estimation.
2.1.1 Image Embeddings
We employ deep learning models that were originally trained for object detection and classification tasks. Specifically, we use four pre-trained convolutional neural networks available in the Keras Deep Learning Library: VGG19, ResNet50, InceptionV3, Xception. [^8] VGG19, a deep convolutional neural network with 19 layers (Simonyan and Zisserman, 2015). ResNet50, is a 50-layer convolutional neural network (He et al., 2016). InceptionV3 and Xception are convolutional neural networks with 48 and 71 layers (Szegedy et al., 2016; Chollet, 2017). All four models achieve high predictive accuracy on the ImageNet validation dataset, and are widely known for their strong performance in image classification tasks. [^9] The ImageNet data (https://www.image-net.org/) is used for the ImageNet Large Scale Visual Recognition Challenge (ILSVRC), which evaluates algorithms for object detection and image classification, a benchmark in object category classification (Russakovsky et al., 2015).
@@ 2.1.1 Image Embeddings
We extract image embeddings from product photos, which researchers can easily obtain via online search or by scraping product pages from e-commerce platforms. Each model transforms the original image into a vector representation – an embedding – and classifies an image from its embedding by predicting the object it contains. These models were originally trained to classify images into labeled categories (e.g., “cup,” “book,” or “sofa”). However, since our goal is not to label products but to measure visual features that distinguish them from each other, we remove the classification layer from these models and work directly with embeddings. Because these models perform well at distinguishing similar objects, we expect the embeddings to capture key visual features that differentiate products.
We do not commit to a single model and intentionally use multiple deep learning models with different architectures. As we explain in Section 3.2, we select the one that best explains substitution patterns in the data.
2.1.2 Text Embeddings
Next, we extract information from text data, including product titles, descriptions, and customer reviews. We use two simple but widely used “count” models as well as pre-trained models with varying levels of sophistication, generating a separate set of embeddings for each model and type of text data. Before extracting embeddings, we pre-process text data as described in Appendix A.
First, we use a bag-of-words count model, which transforms text into fixed-length vectors by counting word occurrences. This model does not consider word order and only measures how often words appear in a document. Second, we apply a bag-of-words model with a TF-IDF vectorizer. Although similar to the first method, TF-IDF assigns more weight to words that are frequent in a specific text but rare across others, thus emphasizing unique words. This makes it more likely that text embeddings capture distinctive words that differentiate products from one another.
Next, we use the Universal Sentence Encoder (USE), a transformer-based language model (Cer et al., 2018), and the Sentence Transformer (ST) model, a pre-trained Sentence-BERT model designed to produce semantically meaningful sentence embeddings (Reimers and Gurevych, 2019). [^10] The model trained by Reimers and Gurevych (2019) is a more efficient version of the widely used BERT network (Devlin et al., 2018). Unlike the bag-of-words models, both USE and ST account for word order and the context in which words appear within sentences. Embeddings from these models are commonly used for text classification, semantic similarity analysis, clustering, and other natural language processing tasks. Both models achieve excellent performance on semantic textual similarity benchmarks (Cer et al., 2017).
While count models are relatively simple, they can detect individual attributes mentioned in titles, descriptions, and reviews. In contrast, the more advanced ST and USE models are better at capturing nuanced similarities in how sellers and consumers describe products. For example, they can identify when sellers or consumers describe the same functional benefits using similar language, even if the words are not exactly the same. Thus, we view count models as a useful benchmark that tells us whether our method merely extracts attributes mentioned in texts or captures more subtle cues about product substitution.
2.2 Generating Principal Components
@@
Embeddings provide a lower-dimensional representation of the original texts and images. However, they remain high-dimensional compared to the number of attributes researchers typically include in demand models. For example, it would be impractical to incorporate 512-dimensional VGG19 embeddings into a random coefficients logit model, as it would require computationally intensive numerical integration over all dimensions. For this reason, we further reduce the dimensionality by applying Principal Component Analysis (PCA) to the embeddings.
For a given category, we extract principal components from each set of embeddings corresponding to a given model and data type (e.g., USE embeddings from customer reviews) and incorporate them into a standard mixed logit model (Berry et al., 1995, 2004), estimating a separate random coefficient for each component. Principal components are particularly appealing because they are orthogonal to each other, which mitigates multicollinearity and simplifies the estimation of random coefficients in the mixed logit model (Backus et al., 2021). We perform model selection by choosing the combination of principal components that minimizes the AIC (see Section 3.2.1).
@@
Because we apply PCA separately within each category, the extracted principal components capture the key dimensions that differentiate products within that category. By contrast, raw embeddings come from models trained on large, general-purpose datasets. For instance, image embeddings come from models trained to classify images into broad categories, such as “tablet,” “laptop,” or “cellphone,” and thus contain features that distinguish images across these categories. However, to estimate demand, we need to analyze consumer choices within a given category, where products often share visual features due to similar designs. PCA helps us filter out the variation in embeddings necessary for sorting products into categories, allowing us to focus on dimensions most relevant for analyzing substitution patterns within each category.
2.3 Interpretation of Principal Components
Standard mixed logit models capture substitution by measuring how similar products are in terms of their observed attributes. Similarly, our approach captures substitution by measuring how similar products are in terms of their visual designs and textual descriptions.
There are several reasons why this proximity may be linked to substitution. Consumers may choose based on how a product looks and how it is described by sellers or other consumers. For instance, a consumer who finds one product visually appealing may also be drawn to similar-looking products, and they might expect products with similar descriptions to meet their needs in a similar way.

### v3

@@
Demand Estimation with Text and Image Datathanks: We thank Steve Berry, Alex Burnap, Matt Gentzkow, Joonhwi Joo, Philip Haile, Yufeng Huang, Jon McClure, Carl Mela, Sanjog Misra, Olivia Natan, Matt Osborne, Katja Seim, Adam Smith, Yuyan Wang, Kevin Williams and seminar participants at Bocconi University, Chicago Booth, London Business School, Stockholm School of Economics, University of Cologne, Nova SBE, Marketing Science Conference 2022, SICS Conference 2024, AI/ML Yale Conference, and SED Conference for helpful comments. We also thank Celina Park, Tanner Parsons, James Ryan, Janani Sekar, Andrew Sharng, Franklin She, Adam Sy, and Vitalii Tubdenov for outstanding research assistance. Contact information: Giovanni Compiani (gio1compiani@gmail.com), Ilya Morozov (ilya.morozov@kellogg.northwestern.edu), Stephan Seiler (stephan.a.seiler@gmail.com). Giovanni Compiani gratefully acknowledges support from the Center for Applied AI at Booth.
@@
We propose a demand estimation method that leverages unstructured data to infer substitution patterns. Using pre-trained deep learning models, we extract embeddings from product images and textual descriptions and incorporate them into a mixed logit demand model. This approach enables demand estimation even when researchers lack data on product attributes or when consumers value hard-to-quantify attributes, such as visual design. Using a choice experiment, we show this approach outperforms standard attribute-based models at counterfactual predictions of second choices. We also apply it to 40 product categories on Amazon.com and consistently find that unstructured data identifies close substitutes within each category.
@@ 1 Introduction
In this paper, we show how researchers can incorporate text and image data in demand estimation to recover substitution patterns. We propose an approach that uses product images, titles, descriptions, and customer reviews—unstructured data that are widely available even in markets where collecting product attributes is challenging. Using pre-trained deep learning models, we transform images and texts into vector representations—embeddings. We then apply Principal Component Analysis (PCA) to reduce dimensionality and capture the main dimensions of product differentiation. We incorporate these principal components into a mixed logit demand model (Berry et al., 1995; Nevo, 2001), interacting them with random coefficients similar to how researchers usually treat observed product attributes in demand estimation.
@@
This approach enables researchers to incorporate hard-to-quantify product attributes, such as visual design from images and functional benefits from text, while also circumventing the need to subjectively choose which observed attributes to collect. Given the prevalence of unstructured data, we see our approach as a valuable addition to the toolbox of empirical researchers. To make it easier for others to apply our approach, we provide a publicly available Python package, DeepLogit, alongside this paper. [^2] The package is available on PyPI and in the public GitHub repository: github.com/deep-logit-demand/deeplogit. Our package heavily borrows from the existing package Xlogit for GPU-accelerated estimation of mixed logit models, developed by Arteaga et al. (2022).
To show that the proposed method performs well at recovering substitution patterns, we validate it extensively on several datasets. Validation in demand estimation is challenging because substitution patterns are usually unobserved, leaving no clear “ground truth” to evaluate model predictions. To overcome this challenge, we design a choice experiment where participants choose a book from a list of options. We randomize both prices and rankings to identify substitution patterns. Crucially, we elicit both first and second choices, estimate different demand models using only first choices, and evaluate their performance based on how well they predict second choices counterfactually. [^3] We refer to second choice predictions as “counterfactual predictions” in the sense that second choices are not used during estimation, which is based solely on first choices. Conditional second choices are informative about substitution because they reveal which product a consumer would substitute to if their preferred option was unavailable. Recording second choices thus enables us to assess how well each demand model predicts substitution patterns. [^4] Second choices are related to the notion of diversion ratios, which reflect the degree of substitutability and thus the intensity of price competition between two products (Shapiro, 1995; Conlon and Mortimer, 2021; Conlon et al., 2023). Recovering them is crucial for antitrust authorities trying to predict price increases from horizontal mergers and for managers optimizing assortments and prices.
Using these experimental data, we show that both text and images contain useful information about substitution patterns. Intuitively, images capture substitution because book covers visually convey information about genre. Text performs even better, likely because descriptions and reviews capture nuanced details about book plots, helping identify similar books even within the same genre. We then show that the best performing model—based on customer reviews—improves counterfactual predictions of second-choices relative to the standard mixed logit with observed attributes. In fact, the improvement relative to that model is on par with the improvement that the mixed logit achieves over the plain logit model without random coefficients. Thus, in this application, we can recover substitution from unstructured data alone, and in fact can achieve better performance than is possible even with attribute data.
To assess the economic importance of these improvements, we simulate hypothetical mergers of book publishers. We find that differences in estimated substitution patterns are large enough such that an antitrust agency relying on the standard mixed logit with attributes might incorrectly approve mergers that should be challenged and block mergers that should be approved. Thus, our results show that incorporating text and image data yields improvements in estimated substitution patterns that are both statistically and economically significant.
We also validate our approach on observational data. We apply it to 40 product categories on Amazon.com, including groceries, pet food, office supplies, beauty products, electronics, video games, and clothing. We combine data on online purchases from the Comscore Web Behavior panel with product images, titles, descriptions, and reviews collected from Amazon’s product detail pages. Here, our contribution is twofold. First, we show that text and images contain useful signals about substitution in all studied categories, highlighting that our approach is broadly applicable across markets. Second, we find—perhaps unexpectedly—that text performs best in some categories where images might be expected to be more useful, such as clothing, and vice versa. A practical takeaway is that researchers may benefit from collecting both text and images, even when only one seems relevant a priori.
Our paper contributes to the extensive literature on demand estimation. Specifically, we build on papers that model substitution through the heterogeneity of preferences over observed product attributes (Berry et al., 1995; McFadden and Train, 2000; Nevo, 2001; Berry et al., 2004). We show that incorporating the principal components of text and image embeddings into these models captures substitution patterns well. A key contribution of our paper is that we validate this approach extensively using both experimental and observational data.
This paper is also related to a vast literature in computer science that transforms unstructured data into lower-dimensional representations (Krizhevsky et al., 2012; Mikolov et al., 2013; Goodfellow et al., 2016). There is a growing literature on applications of pre-trained deep learning models in economics and marketing (see Battaglia et al., 2024 for a review). While most prior work focuses on predicting observed outcomes, our goal is fundamentally different: we aim to predict counterfactual quantities, such as substitution patterns. Predicting these quantities well is crucial for applications like merger simulations and optimal pricing. Yet, they are usually unobserved in choice data. As a result, we cannot apply standard cross-validation techniques and instead need to design a custom experiment that allows us to measure these otherwise unobserved quantities. To our knowledge, we are the first to use such an experiment for validation of demand models. [^5] Berry et al. (2004) show that second choice data are useful for estimating substitution patterns, and Conlon et al. (2022) estimate demand from aggregate data on first-choice probabilities and a subset of second-choice probabilities. Conceptually, our validation approach reverses the logic of these papers: we only use first-choice data for estimation and treat second choices as “holdout” data for model validation. Additionally, Raval et al. (2022) use hospital closures induced by natural disasters to compare the ability of various models to predict diversions.
@@
Several other papers have proposed leveraging text and image data in demand estimation. [^6] Netzer et al. (2012) use data on the co-occurrence of products mentioned in online discussion forums to generate a visual representation of competing products, but they do not incorporate these measures into demand estimation. Dew (2024) uses image embeddings from pre-trained models to predict consumers’ ratings for products based on past ratings elicited in a survey. Sisodia et al. (2024) extract interpretable product attributes from images and use them to estimate preferences. Our approach is simpler and scales better across product categories, and thus, it may be preferable in applications where interpretability of extracted attributes is not essential. Quan and Williams (2019) incorporate product image embeddings as shifters of mean utilities. Similarly, Lee (2024) employs large-language models to predict utility intercepts of new products based on their textual descriptions. These methods predict mean utilities but are not designed to estimate utility covariances, which play a crucial role in shaping substitution. [^7] For example, Lee (2024) inverts a plain logit model without random coefficients to extract product fixed effects for the prediction part of his analysis. Thus, his model restricts substitution patterns due to the IIA property. In contrast, our approach models utility covariances and can thus capture flexible substitution patterns—a key input for many empirical applications of interest. Further, while utility intercepts can be recovered from market shares, our goal is to capture substitution patterns that are unobserved. We measure them in an experiment by eliciting second choices, and we use these data for validation. [^8] A recent paper by Han and Lee (2025) uses unstructured data to estimate substitution patterns among fonts in order to study the effects of copyright protection policies.
@@
A possible alternative to our use of unstructured data is to estimate substitution patterns using survey data. For example, Dotson et al. (2019) ask survey participants to rate each product image and then incorporate rating correlations as shifters of utility correlations into a demand model. Similarly, Magnolfi et al. (2022) elicit relative rankings from survey participants (e.g., “Product A is closer to B than to C”), generate embeddings from these rankings, and use them in a random coefficients logit model. Our approach complements these survey-based methods but has the advantage of using only widely available text and image data, thereby avoiding the need for costly category-specific surveys. Another complementary approach is proposed by McClure (2025), who shows how publicly available recommendations data from online platforms can also be used to estimate substitution patterns.
@@
Our approach is well-suited for several empirical applications. Researchers can use it to address a wide range of economics and marketing questions that require accurate and flexible estimates of product substitution. This includes analyzing how horizontal mergers (Nevo, 2000; Federal Trade Commission, 2022), new product launches (Hausman, 1994; Petrin, 2002), corrective taxes (Allcott et al., 2019; Seiler et al., 2021), and trade restrictions (Goldberg, 1995; Berry et al., 1999) influence consumers’ choices and welfare through altering assortments and prices. Additionally, researchers can apply this approach to study how multi-product firms optimize prices and promotions (Hoch et al., 1995; Vilcassim and Chintagunta, 1995). Finally, because our approach circumvents the need to collect distinct attributes for each category, it may be valuable for researchers trying to estimate demand across many categories (Döpper et al., 2024). [^9] Döpper et al. (2024) remark that estimating demand across multiple categories “would be difficult to implement at scale because it would require category by-category assessments about which characteristics are appropriate to include and whether or not relevant data are available.” Our approach can address this problem if researchers have access to unstructured data.
@@ 2 Proposed Approach
The proposed approach involves three steps: (1) extracting embeddings from images and texts, (2) reducing the dimensionality of these embeddings using PCA, and (3) including the resulting principal components into a standard attribute-based logit model with random coefficients.
Step 1. Extracting Embeddings from Texts and Images
@@
To extract information from product images, we extract their low-dimensional representations—embeddings—using pre-trained deep learning models. We use four convolutional neural networks: VGG19 (Simonyan and Zisserman, 2015), ResNet50 (He et al., 2016), InceptionV3 (Szegedy et al., 2016), and Xception (Chollet, 2017). [^10] These models were originally trained to classify images into labeled categories (e.g., “cup,” “book,” or “sofa”). However, since our goal is not to label products but to measure visual features that distinguish them from each other, we remove the classification layer from these models and work directly with embeddings. The key advantage of using pre-trained models is that it reduces the computational burden of estimation while leveraging models trained on large-scale datasets, known for their strong performance in image classification tasks (Russakovsky et al., 2015). Because these models perform well at distinguishing visually similar objects, we expect embeddings to capture the key visual features that differentiate products. Rather than committing to a single model, we perform model selection to determine which one best captures substitution patterns, as described below.
We also extract information from texts, including product titles, descriptions, and customer reviews. We use a bag-of-words model, which represents text as fixed-length vectors based on word counts, as well as a variation with a TF-IDF vectorizer. [^11] TF-IDF assigns more weight to words that are frequent in a specific text but rare across others, thus emphasizing unique words. This makes it more likely that text embeddings capture distinctive words that differentiate products from one another. While these count models are relatively simple, we view them as useful benchmarks because they can still detect attributes mentioned in titles, descriptions, and reviews. In addition, we employ two pre-trained deep learning models: Universal Sentence Encoder (USE) (Cer et al., 2018), and BERT Sentence Transformer (ST) (Reimers and Gurevych, 2019). [^12] The model trained by Reimers and Gurevych (2019) is a more efficient version of the widely used BERT network (Devlin et al., 2018). Both USE and ST produce semantically meaningful sentence embeddings and achieve excellent performance on semantic textual similarity benchmarks (Cer et al., 2017). As a result, they can identify when sellers or consumers describe the same product attributes and functional benefits using similar language, even if the exact words differ.
Step 2. Generating Principal Components
@@
Although embeddings compress texts and images into lower-dimensional representations, they remain high-dimensional compared to the number of attributes typically used in demand models. For example, incorporating 512-dimensional VGG19 embeddings into a random coefficients logit model would require prohibitively costly numerical integration over all dimensions. We therefore apply PCA to further reduce dimensionality (Backus et al., 2021).
@@
The dimensionality reduction via PCA is attractive for multiple reasons. Raw embeddings are trained on general-purpose datasets to classify images into broad categories, such as “tablet” or “laptop.” However, to estimate demand, we need to analyze consumer choices within a given category. PCA helps us filter out the variation in embeddings necessary for sorting products into categories, allowing us to focus on dimensions most relevant for analyzing substitution within a category. Additionally, principal components are appealing because they are orthogonal to each other, which avoids multicollinearity issues and simplifies estimation of random coefficients in demand models.
Step 3. Including Principal Components into a Demand Model
We include principal components into a standard mixed logit model (Berry et al., 1995, 2004), estimating a separate random coefficient for each included component. Since each deep learning model and data type produces different principal components, a key question is which to include in the demand model. [^13] While we could select the specification that best matches observed second choices in our experiment, this strategy is unavailable to most researchers who usually only have first-choice data. We thus propose a model selection algorithm that relies solely on first choices. We perform model selection using in-sample $AIC$. An advantage of using $AIC$ is that it can be interpreted as a bias-corrected estimator of the expected relative Kullback-Leibler (KL) information based on the maximized log-likelihood function (Akaike, 1998). Further, in our experimental results, in-sample $AIC$ strongly correlates with counterfactual performance, justifying its use for model selection (see Section 3.3 where we also implement $BIC$ and cross-validation).
Our model selection algorithm considers a pool of candidate variables consisting of price and the first $P$ principal components. In our application, we choose $P=6$, which collectively account for 70-80% of variance in embeddings (see Appendix Figure A4). We first estimate all possible models with a random coefficient on a single variable and retain the best-fitting model. We then estimate all possible models with two random coefficients and assess whether fit improves relative to the best single-variable model. We continue to increase the number of variables $K$ that have random coefficients until there is no further $AIC$ improvement. We repeat this across all specifications—defined as a combination of a text or image model and data type (e.g., USE Reviews)—and choose the one with the lowest overall $AIC$. The full procedure is given in Algorithm 1.
This algorithm avoids costly combinatorial search over all possible specifications. At the same time, it ensures that we do not rely too heavily on the ordering of principal components, which reflects the explained variance of embeddings but need not be predictive of preferences. For example, the approach might select a model that has random coefficients on the second and fourth principal components, but not the first or third.
      Algorithm 1: Model Selection
Input: - Products $j=1,\dots,J$ - Specifications $m=1,\dots,M$ (all combinations of unstructured data and pre-trained text or image models) - Maximum number of principal components $P$ (researcher’s choice) - Demand model with utility: $u_{ij}=\delta_{j}+\alpha_{i}\text{price}_{ij}+\theta_{i}^{\prime}PC_{j}+\varepsilon_{ij}$ ($\delta_{j}$ = product fixed effects)
1: For each specification $m$ do
2:   Extract embeddings $e^{(m)}=(e_{1}^{(m)},\dots,e_{J}^{(m)})$
3:   Extract principal components $PC_{1}^{(m)},\dots,PC_{P}^{(m)}$ from embeddings $e^{(m)}$ using PCA
4:   Initialize model selection at $K\leftarrow 0$ (zero random coefficients)
5:   Set $\text{BestSpec}^{(m)}$ as Plain Logit and $\text{BestAIC}^{(m)}$ as the corresponding AICAIC
6:   while $\text{BestAIC}^{(m)}$ decreases do
7:     $K\leftarrow K+1$
8:     Estimate mixed logit models with all possible combinations of random coefficients
     on subsets of $K$ variables from the candidate set $\mathcal{R}=\{\text{price, P$C_{1}^{(m)}$, $\dots$, P$C_{P}^{(m)}$}\}$
9:     Find subset $R_{K}^{*}\in\mathcal{R}$ that minimizes $AIC$ at $AIC_{K}^{*}$
10:     if $AIC_{K}^{*}<\text{BestAIC}^{(m)}$ do
11:       $\text{BestSpec}^{(m)}\leftarrow R_{K}^{*}$ update the best specification
12:       $\text{BestAIC}^{(m)}\leftarrow AIC_{K}^{*}$ update the lowest AICAIC
13:   end while
14: end for
15: Choose the best-fitting specification $m^{*}=\arg\min_{m}\text{BestAIC}^{(m)}$
Why Does This Approach Work?